if a+1=b*2 and b+1=a what is a?
Answer by stanbon(75887)

a+1=b*2
b+1=a
----------------
what is a?
b = a-1
Substitute to solve for "a":
a+1 = (a-1)^2
a+1 = a^2 -2a + 1
a^2 -3a = 0
a(a-3) = 0
a = 0 or a = 3
===================
Cheers,
Stan H.